OpenStudy (anonymous):

What is the product in simplest form? State any restrictions on the variable. z^2/z + 1 times z^2 + 3z+2/z^2 + 3z

OpenStudy (anonymous):

\[\frac{ z^2 }{ z + 1}\times \frac{ z^2 + 3z + 2 }{ z^2 + 3z }\]

OpenStudy (anonymous):

@tcarroll010 how do you do this one ?

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

@cgreenwade2000 do you know how to do this !?

OpenStudy (cgreenwade2000):

start by breaking down (x^2 + 3z + 2)

OpenStudy (cgreenwade2000):

It's easier actually to just factor everything as much as possible and then reducing from there.

OpenStudy (anonymous):

(x + 2)(x + 1) ?

OpenStudy (cgreenwade2000):

Yes. Now factor z^2 and z^2+3z

OpenStudy (anonymous):

um idk that

OpenStudy (anonymous):

@cgreenwade2000

OpenStudy (cgreenwade2000):

z^2 = z*?

OpenStudy (cgreenwade2000):

and z^2 +3z = ? There is something you can pull from those terms

OpenStudy (anonymous):

z ?

OpenStudy (cgreenwade2000):

yes so you can simplify z^2 to get z*z what can you factor out of z^2 + 3z?

OpenStudy (anonymous):

z(z +3) @cgreenwade2000

OpenStudy (cgreenwade2000):

Yes. Now rewrite your problem with your factored expressions in

OpenStudy (anonymous):

z*z x (x + 2)(x + 1)/z(z +3) ? @cgreenwade2000

OpenStudy (cgreenwade2000):

 You put some x's in there. Be careful of that. There are no x variables in this problem